OpenStudy (anonymous):

What are the factors of 5x2 − x − 6?

OpenStudy (anonymous):

What are the factors of 5*6=30 that will give you -1?

OpenStudy (anonymous):

Is (5x − 1)(x + 6) the answer?

OpenStudy (anonymous):

Do you understand how you got that answer?

OpenStudy (anonymous):

I saw you put -1 and that was the only answer choice with an -1. Lol

OpenStudy (anonymous):

Oh... hmm, well to figure this out if you have to show working is you multiply the coefficient of the x^2 which is 5 by the constant which is -6 to give you -30 and factors of -30 are 1, 2, 3, 5, 6, 10, 15 and 30

OpenStudy (anonymous):

@purplec16 It's quite strange to ask ''What are the factors of 5*6=30 that will give you -1'' Since it's NOT 5*6 in this case You may factor it in this way, 5x^2 − x − 6 = 5x^2 +5x - 6x -6 = 5x (x+1) -6(x+1) = (5x-6)(x+1)

OpenStudy (anonymous):

So then you say well what two factors can give me -1 and from the factors it looks like it is -6x and 5x: =5x^2-6x+5x+6 =5x^2+5x-6x+6 =5(x+1)-6(x+1) =(5x-6)(x+1)

OpenStudy (anonymous):

Hey guys, please look at the question carefully... It is 5x^2 − x − 6!!! The constant term is -6, not -30 or 30!

OpenStudy (anonymous):

I got it guys, thanks so much (:

OpenStudy (anonymous):

You're welcome... and I do apologize if I may have confused you so please take it to consideration the corrections that they made in using the correct terms so you don't confuse yourself, lol... like I may have confused you.

OpenStudy (anonymous):

Lol, math in general confuses me.